OpenStudy (anonymous):

FInd the length of an arc with a radius 2 inches and a central angle of 45

jimthompson5910 (jim_thompson5910):

s = (angle/360)*2*pi*r s = (45/360)*2*3.14*2 s = ???

OpenStudy (anonymous):

1.570...

jimthompson5910 (jim_thompson5910):

yes good

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw